# MedPhoton GmbH

### v1

@@ 1 Introduction
Segmentation is a crucial field of automated medical image analysis, for a multitude of applications, from segmenting low-level (e.g., nuclei) to high-level tissue structure (differentiation between cancer, stroma, and necrosis). Classical segmentation methods (thresholding, region-based, watershed), though useful for both 2D and 3D segmentation, often struggle with the complexity of pathological and anatomical structures, particularly high-level structures observed in medical imaging. Deep learning methods exhibit the state-of-the-art for semantic segmentation, commonly employing encoder-decoder models [1, 2, 3]. These models are trained to encode latent information in a bottleneck and then use this information to reconstruct a mask (segmentation) or the original volume (image reconstruction). Fully-convolutional networks, exemplified by Unet [4], utilize this encoder-decoder structure to generate segmentation masks. Models like nn-unet [1] extend Unet’s architecture to 3D and demonstrate state-of-the-art performance on various computed tomography (CT) segmentation tasks [5]. UnetR [2] and SegFormer [3] adapt nn-unet by incorporating transformer blocks like multihead attention.
Multi-task learning aims to share knowledge over multiple tasks of a common model for knowledge transfer, regularization and data efficiency. E.g. multi-task learning approaches have previously been investigated by Weninger et al. [6] to reduce the amount of needed labelled training data by adding reconstruction to a segmentation task. Similarly, Mlynarski et al. [7] include a task of 2D tumor detection score estimation. Both approaches use encoder-decoder structures. Therefore, the addition of multiple tasks increase the amount of data that can be used to train their encoder.
We propose a novel, model agnostic, extension to semantic segmentation. By formulating model optimisation through multi-task learning of segmentation and image reconstruction we facilitate morphology based regularisation. For experimental validation, we select the state-of-the-art for automated semantic segmentation, 3D nn-unet [1, 5]. The evaluation is conducted using the well-established Liver Tumor Segmentation Benchmark (LiTS) [5]. To assess our proposed adaptation at different data quality, we synthesise cone-beam CT (CBCT) from the original LiTS, with varying numbers of used projections employed for CT reconstruction. We further evaluate the effect of using the CBCT corresponding to the highest quality volumes as the reconstruction target, facilitating a denoising effect.
@@ 2 Method
As shown in Fig. 1, a 3d nn-unet [4, 1] is chosen as the baseline (red path). This unet consists of an encoder with $3$ double convolution layers and $3\times 3\times 3$ convolutional kernels, connected by 3d max pooling. The latent space consists of one double convolution block leading into the unet decoder which mirrors the encoder. As is typicall for unet, each double convolutional output in the encoder is also connected to the decoder double convolutional block of the same order. Additionally, one 3d convolutional layer is added to the decoder with a filter size of $1\times 1\times 1$ and the number of filters set to the amount of segmentation classes. The rest of the convolutional layers are set to $\{32,64,128,256\}$ filters along the stream direction in the encoder. Batch norm is applied after each layer in the double convolutional blocks. The model is trained utilizing a sum ($Loss_{1}$) of BCE and Dice $Loss_{1}=BCE(\hat{s},s)+Dice(\hat{s},s)$.
@@ 2.1 Dataset
Evaluation is performed using the LiTS dataset. LiTS consists of $131$ abdominal CT scans in the training set and $70$ test volumes. The $131$ train volumes include segmentations of 1) the liver and 2) liver tumors. The dataset contains data from $7$ different institutions with a diverse set of liver tumor diseases. For further information we refer to Bilic et al. [5]. To evaluate effects on vastly different volume qualities, the LiTS dataset was converted into synthetic CBCT scans. Corresponding projections were simulated per CT volume, with the number of projections $n_{p}\in\{490,256,128,64,32\}$. $490$ projections was chosen as the default configuration to show a visual quality similar to the original CT and $32$ as the lowest configuration, showing little information and many artifacts.
@@ 2.2 Experimental Setup
We evaluate the introduced methods holistically and patch based on liver and liver tumor segmentation [5, 8, 9, 10]. The segmentations are evaluated using the Dice score. A threshold is applied to the model output before Dice calculation to establish binary labels. This threshold is set to $0.5$ in all models and setups. To fit the holistic data on an NVIDIA RTX A6000 grafics card the volumes are downscaled by the factor $2$. For patched segmentation patches of the size $192\times 192\times 192$ are extracted from the full-scale CBCT data. On inference, the patch segmentations are re-agregated before calculating a holistic Dice.
Models and setups are evaluated using the different quality CBCT datasets with the quality depending on the amount of used projections for simulation $n_{p}\in\{490,256,128,64,32\}$. A further investigated parameter is the reconstruction target with the two investigate setups, mt-c and mt-b. The setup mc-c uses the input volume of the current quality as its reconstruction target while mt-b always uses the best possible quality input volume ($v_{o}$) as its reconstruction target. Therefore, in the case $n_{p}=490$ both mt-b and mt-c lead to the same results. All setups are trained and evaluated $4$ times to facilitate stable results with the same random splits for results comparability. Since annotations are not available for the LiTS test dataset, the test dataset was disregarded for this publication and the original train set was separated into training-validation-testing data [8, 9]. The separation was performed using the ratio 0.7, 0.2 and 0.1 ($training:0.7$, $validation:0.2$, $testing:0.1$).

### v2

@@ 1 Introduction
Segmentation is a crucial field of automated medical image analysis, for a multitude of applications, from segmenting low-level (e.g., nuclei) to high-level tissue structure (differentiation between cancer, stroma, and necrosis). Classical segmentation methods (thresholding, region-based, watershed), though useful for both 2D and 3D segmentation, often struggle with the complexity of pathological and anatomical structures, particularly high-level structures observed in medical imaging. Deep learning methods exhibit the state-of-the-art for semantic segmentation, commonly employing encoder-decoder models [isensee2019nnu, hatamizadeh2022unetr, xie2021segformer]. These models are trained to encode latent information in a bottleneck and then use this information to reconstruct a mask (segmentation) or the original volume (image reconstruction). Fully-convolutional networks, exemplified by Unet [ronneberger2015u], utilize this encoder-decoder structure to generate segmentation masks. Models like nn-unet [isensee2019nnu] extend Unet’s architecture to 3D and demonstrate state-of-the-art performance on various computed tomography (CT) segmentation tasks [bilic2023liver]. UnetR [hatamizadeh2022unetr] and SegFormer [xie2021segformer] adapt nn-unet by incorporating transformer blocks like multihead attention.
Multi-task learning aims to share knowledge over multiple tasks of a common model for knowledge transfer, regularization and data efficiency. E.g. multi-task learning approaches have previously been investigated by Weninger et al. [weninger2020multi] to reduce the amount of needed labelled training data by adding reconstruction to a segmentation task. Similarly, Mlynarski et al. [mlynarski2019deep] include a task of 2D tumor detection score estimation. Both approaches use encoder-decoder structures. Therefore, the addition of multiple tasks increase the amount of data that can be used to train their encoder.
We propose a novel, model agnostic, extension to semantic segmentation. By formulating model optimisation through multi-task learning of segmentation and image reconstruction we facilitate morphology based regularisation. For experimental validation, we select the state-of-the-art for automated semantic segmentation, 3D nn-unet [isensee2019nnu, bilic2023liver]. The evaluation is conducted using the well-established Liver Tumor Segmentation Benchmark (LiTS) [bilic2023liver]. To assess our proposed adaptation at different data quality, we synthesise cone-beam CT (CBCT) from the original LiTS, with varying numbers of used projections employed for CT reconstruction. We further evaluate the effect of using the CBCT corresponding to the highest quality volumes as the reconstruction target, facilitating a denoising effect.
@@ 2 Method
As shown in Fig. 1, a 3d nn-unet [ronneberger2015u, isensee2019nnu] is chosen as the baseline (red path). This unet consists of an encoder with $3$ double convolution layers and $3\times 3\times 3$ convolutional kernels, connected by 3d max pooling. The latent space consists of one double convolution block leading into the unet decoder which mirrors the encoder. As is typicall for unet, each double convolutional output in the encoder is also connected to the decoder double convolutional block of the same order. Additionally, one 3d convolutional layer is added to the decoder with a filter size of $1\times 1\times 1$ and the number of filters set to the amount of segmentation classes. The rest of the convolutional layers are set to $\{32,64,128,256\}$ filters along the stream direction in the encoder. Batch norm is applied after each layer in the double convolutional blocks. The model is trained utilizing a sum ($Loss_{1}$) of BCE and Dice $Loss_{1}=BCE(\hat{s},s)+Dice(\hat{s},s)$.
@@ 2.1 Dataset
Evaluation is performed using the LiTS dataset. LiTS consists of $131$ abdominal CT scans in the training set and $70$ test volumes. The $131$ train volumes include segmentations of 1) the liver and 2) liver tumors. The dataset contains data from $7$ different institutions with a diverse set of liver tumor diseases. For further information we refer to Bilic et al. [bilic2023liver]. To evaluate effects on vastly different volume qualities, the LiTS dataset was converted into synthetic CBCT scans. Corresponding projections were simulated per CT volume, with the number of projections $n_{p}\in\{490,256,128,64,32\}$. $490$ projections was chosen as the default configuration to show a visual quality similar to the original CT and $32$ as the lowest configuration, showing little information and many artifacts.
@@ 2.2 Experimental Setup
We evaluate the introduced methods holistically and patch based on liver and liver tumor segmentation [bilic2023liver, araujo2022liver, han2022effective, wang2022automatic]. The segmentations are evaluated using the Dice score. A threshold is applied to the model output before Dice calculation to establish binary labels. This threshold is set to $0.5$ in all models and setups. To fit the holistic data on an NVIDIA RTX A6000 grafics card the volumes are downscaled by the factor $2$. For patched segmentation patches of the size $192\times 192\times 192$ are extracted from the full-scale CBCT data. On inference, the patch segmentations are re-agregated before calculating a holistic Dice.
Models and setups are evaluated using the different quality CBCT datasets with the quality depending on the amount of used projections for simulation $n_{p}\in\{490,256,128,64,32\}$. A further investigated parameter is the reconstruction target with the two investigate setups, mt-c and mt-b. The setup mc-c uses the input volume of the current quality as its reconstruction target while mt-b always uses the best possible quality input volume ($v_{o}$) as its reconstruction target. Therefore, in the case $n_{p}=490$ both mt-b and mt-c lead to the same results. All setups are trained and evaluated $4$ times to facilitate stable results with the same random splits for results comparability. Since annotations are not available for the LiTS test dataset, the test dataset was disregarded for this publication and the original train set was separated into training-validation-testing data [araujo2022liver, han2022effective]. The separation was performed using the ratio 0.7, 0.2 and 0.1 ($training:0.7$, $validation:0.2$, $testing:0.1$).